What is the simplify form of (6^(1/2)+i)(6^(1/2)-i).
i is the imaginary number -1^(1/2)

Answer by Alan3354(69443)

What is the simplify form of (6^(1/2)+i)(6^(1/2)-i).
---------
(a + bi)*(a - bi) = a^2 + b^2
--------
(6^(1/2)+i)(6^(1/2)-i) = 6 + 1 = 7